23. Assume the following information: You have $900,000 to invest: Current spot rate of Australian dollar (AS) 180-day forward rate of the Australian dollar 180-day interest rate in the U.S. 180-day interest rate in Australia $.62 -.64 3.5% 3.0% If you conduct covered interest arbitrage, what is the dollar profit you will have realized after 180 days?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard UO37 UL Font Alignment Number Styles Cells Edting UM UN UO UP UQ UR US UT UV UW UX UY UZ 24 25 26 27 28 29 30 31 32 COVERED INTEREST ARBITRAGE STEP 1 SPOT RATE 1 A$ = 0.625 THEREFORE 900000 DOLLARS = 900000/0.62 = 1451612.90 A$ STEP 2 INVEST IN AUSTRALIA AT 3% FOR 180 DAYS- 1495161.29 UT28*(1.03) STEP 3 CONVERT IN DOLLAR = 956903.23 UT30*0.64 FORWARD RATE = 1 AS = 0.64 34 35 36 37 38 39 40 STEP 4 DOLLAR PROFIT 56903.23 (956903.23 -900000) 11 1 r ri ! KE CAPM UTILITY, SH beta bond c future INDEX | INTL CAP BUD SING FV, ANNUITY DIR cleanYIELD bond stru WACCRESI ex d erences: 130 130% 03:35 01-02-2019